OpenStudy (honeymonster):

how do i make x the subject of the formula with this equation. ax+by+c =0

sam (.sam.):

\[\begin{array}{l} a x+b y+c=0 \\ \text{Subtract }b y+c\text{ from both sides:} \\ a x=-b y-c \\ \text{Divide both sides by }a: \\ x=\frac{-c-b y}{a} \\ \end{array}\]